# Unsourced Random Access: A Comprehensive Survey

### v2

@@ V-D Existing URA Solutions on Alternative Setups
Most URA schemes in the literature consider direct links between all the users and the BS; however, in certain environments, the direct link between some users and the BS may be blocked or significantly attenuated. In these cases, one solution can be employing RIS, a promising technology that can furnish the URA system with high spectral efficiency and energy savings. Specifically, it is shown that a passive RIS equipped with many low-cost passive elements, which can intelligently tune the phase-shift of the incident electromagnetic waves, and reflect them in a desired direction without any amplification, can improve the efficiency of a URA system by enabling line-of-sight paths between the transmitters and the receivers in problematic environments with many blocking obstacles [121, 122]. More specifically, the RIS-based schemes in [123, 124] operate in two phases. In the first phase, called the RIS configuration phase, the BS detects the active pilots and estimates their respective channel coefficients, using which the BS suitably selects the phase shifts of the RIS elements. In the second phase, called the data phase, transmitted messages of active users are decoded using SIC-based decoders. The RIS-based URA solution in [123] outperforms the one in [124], with a power saving of up to 10 dBm. This improvement is due to the employment of a slotted transmission structure, which reduces the interference, as well as a more suitable metric in the RIS design in [123], where the overall phase-shift matrix is obtained by minimizing the decoding error probability, in contrast to maximizing the minimum channel gain among active users for the RIS design in [124]. The authors improve the RIS design metrics of [123] in [125] and obtain some bounds on its performance.
@@
Conventionally, in MIMO URA systems, the ideal hardware units are assumed at both the BS and the UE side. However, due to the massive number of users and the antenna elements, the hardware needs to be inexpensive, and these elements are prone to hardware impairments (HWI) such as I/Q imbalance, phase noise, and quantization errors. The problem of URA with HWI is considered in [126], where the solution developed in [14] is adapted to the presence of HWI.
@@
While most URA works consider the transmission over i.i.d. Rayleigh block-fading channels, some investigate other channel models. A channel model with one single dominant path between a user and the BS is considered in [127], and an approach similar to the one in [114] is proposed, exploiting the user location geometry by linking the transmission features such as the employed preamble or interleaver to the angle of arrival of the user. The dynamics of the user states and channels are taken into account in a slotted structure in [128], namely, the users can start or end their transmission in any slot and they can be active in a subset of the slots, and the channel coefficients change from one slot to another. The authors design a two-step decoding algorithm with an inner AMP and an outer variational message passing. Type-based MA, where each device observes the state of a process and transmits it to a common receiver is investigated in an unsourced setup in [129], where the receiver’s goal is to estimate the states and their multiplicity. The problem of user identification and authentication is applied to URA in [130] with the aid of a message authentication code. Moreover, the authors of [131] propose a URA scheme combining physical and medium access layer approaches where medium access layer slotting carries information. They show that carrying information by the medium access control layer improves the system performance. The Rician fading channel is considered in [132] through an application of coded compressed sensing. A correlated fading scenario is tackled in [133]. Frequency-selective fading is examined in the context of MIMO URA in [134] and [135]. The authors of [134] propose a tensor-based modulation scheme, while a slotted scheme employing OFDM is studied in [135].
@@
The asynchronous setup receives increasing attention due to its easier implementation in hardware, and it is studied in MIMO URA in [136, 138, 137]. A slotted concatenated coding scheme is proposed in [136] based on the SPARCs and tree codes in an OFDM system. The authors of [137] tackle both the timing and frequency offsets by proposing an asynchronous URA scheme by employing SPARC, LDPC codes, and multistage orthogonal pilots. Moreover, symbol asynchronicity in URA is investigated in [138].
@@
Furthermore, some works have investigated URA in other communication setups such as cell-free MIMO [139, 140, 141] and erasure [142, 143] channels. For instance, a cell-free setup is considered in [139] where the coverage area is divided into multiple regions each served by a receiver unit. The users in each region employ a common codebook different from other regions to transmit their messages, and the received signals at the receiver units are jointly processed in a central unit. Joint processing is also considered in [140] and [141], where a combination of channel coding and fixed sparse coding is employed in the former, while an uncoupled CCS scheme utilizing coarse location information for trellis stitching is proposed in the latter.
@@
The common assumption in the URA literature is that all the users have the same number of bits to transmit or require the same quality of service. However, this may be impractical in situations where the devices have different payload requirements or power budgets, especially considering their massive number. This problem, which is called multi-class URA is addressed in [144, 145] in the coded compressed sensing framework. More specifically, a scenario is considered where devices are clustered into two classes with different SNR levels and payload requirements. In the cluster with higher power, devices transmit using a two-layer superposition modulation, while those in the other cluster transmit with the same base constellation as in the first cluster. In order to recover packets at the receiver, a CCS-based approach is employed, where an outer code is used to stitch message fragments together across times and layers.
Now that we have reviewed the existing URA solutions, we provide a summary of different URA setups and the corresponding solution approaches in Fig. 20.
Figure 20: A summary of different URA setups in the literature and their corresponding common approaches.
VI Remarks and Future Research Directions
In this survey, we had a comprehensive look at the URA literature with an emphasis on the state-of-the-art. We grouped the different URA solutions and explained the representative work in each in more detail while clarifying its connection to related works. Moreover, we compared the works in the literature both setup-wise and performance-wise in order to have a clearer view of the existing URA solution.
@@ VI Remarks and Future Research Directions
As evident from this survey, there have been numerous works on URA in recent years. However, there are still some issues, some critical, that need to be investigated further. We will next outline some of these issues and provide several future research directions. Moreover, we highlight potential methods of addressing them.
VI-A Asynchronous URA
@@
In URA, the common assumption is that the users transmit simultaneously; or more precisely, their signals are received at the same time at the receiver (e.g., symbol synchronicity). However, due to the nature of the users and devices in envisioned URA applications such as massive sensor networks, such transmission coordination is very challenging. This makes asynchronous URA an important research direction to investigate. Note that even if the transmission synchronization is realized, the synchronous reception cannot be guaranteed. The reason is that due to the massive number of potential devices, and hence diverse distances to the receiver, the difference in reception times between many users can be larger than a time slot, which makes the setup asynchronous. It is shown that if the asynchronicity is dealt with properly, it can lead to performance improvement even compared to the synchronous case [146].
@@
There have been some attempts to address the asynchronous setup [91, 90, 95, 136, 137, 51], however, with a bounded time horizon, which enable them to convert the problem to one similar to the synchronous case, e.g., by employing a limited decoding time frame. For instance, in [91], the authors employ a frequency-domain solution to convert the asynchronous problem to a phase-shift estimation one. Then, conventional synchronous strategies such as TIN-SIC are utilized to decode the active users.
The only work that focuses on a fully asynchronous (i.e., with an unbounded time horizon) URA setup is [147], where the authors consider a window-based decoding for GMAC URA. In this scheme, each user transmits a preamble (from a pool of available ones) for timing acquisition concatenated with a channel-coded payload, and the receiver uses a SIC-based sliding window for recovering the transmitted packets.
@@
URA-specific challenges: The main URA-specific challenge of a fully asynchronous setup for fading URA is the computational complexity, which is much higher than the conventional (non-URA) setup due to the higher number of required preambles. More specifically, in an asynchronous GMAC setup, the preambles only serve to identify the start of the packets at the receiver, so a preamble pool of even a very small size would suffice; while in an asynchronous fading setup, they are also needed to estimate the channel coefficients, which are then used for user separation. In a conventional fading setup, the receiver only searches over the assigned preambles, whose number is equal to the number of active users. However, in a URA setup, due to the lack of coordination between the users and the receiver, the number of available preambles has to be much higher than in the conventional setup, hence much higher computational complexity for user activity detection and channel estimation. This computational complexity significantly increases in the asynchronous setup with a sliding window receiver, as the receiver shifts and correlates all possible preambles through the window to find the start of the packets as well as the user channel estimates, repeating this process with each shift of the sliding window. This highly increases the overall computational complexity, which can even make the solution practically infeasible. One way to reduce the complexity is to increase the moving steps of the sliding window, at the cost of lower overall performance. However, smarter schemes are needed to alleviate this issue.
VI-B Robust URA
@@
Like any other communication system, the performance of a URA system is affected by the presence of imperfections in the assumed models or employed communication blocks. For instance, the presence of hardware impairments where the employed hardware is not as ideal as assumed, e.g., due to non-linearities of the power amplifiers or phase errors in phase-shifters, which is very probable in many URA applications where the devices are actually cheap and low-precision sensors. Only one work in the URA literature has addressed this issue [126], where a robust solution is presented for the case of residual hardware impairments in which the remaining hardware impairments after compensation are statistically modeled as additive and multiplicative noises. Therefore, the analysis of different hardware impairments on the performance of existing URA solutions and providing robust solutions for them is an interesting research direction.
@@
URA-specific challenges: Looking at the URA literature, we can observe that the URA schemes are more sensitive than the conventional ones to the choice of design parameters. That is, in order for the URA schemes to have acceptable performance, their design parameters need to be optimized every time a system hyper-parameter, such as active user load or the target PUPE, changes. This makes their standardization and practical employment challenging. Therefore, modifications are needed to make the URA solutions robust against changes in system hyper-parameters. One way is to employ stochastic optimization techniques, e.g., optimizing the expected value of the cost metric (e.g., PUPE) over the statistics of the intended system hyper-parameters in order to obtain the set of design parameters that work well in a wide range of system hyper-parameters.
VI-C Machine Learning-Aided URA
@@
Current solutions look at the URA problem in a single shot. However, the transmission is continuous and spans several frames in many practical applications. In these cases, machine learning (ML) techniques such as reinforcement learning can be used to adapt to the environment and learn suitable transmission strategies. ML techniques can also be used to simplify some blocks in the decoding process, such as user activity detection (similar to the deep learning based solutions in [148]) and channel estimation. For instance, in [149], the authors show that employing a dictionary learning based solution obviates the need for pilots for channel estimation in the fading URA scenarios while maintaining an acceptable performance. URA-specific challenges: Apart from general challenges in implementing ML solutions in communication systems, such as energy and computational constraints and compatibility with legacy systems, there are some that are specific or more pronounced in a URA setup, which is mostly due to the massive number of potential users (up to the order of millions) and their sporadic activity. Training a neural network with an input of very large size (due to the massive number of potential users) with non-stationary characteristics (due to the sporadic activity) is a very challenging task.
It can be vice versa as well: URA techniques can be employed to better realize some ML-based solutions. For instance, in federated learning (FL), the receiver is only interested in the sum of the transmitted signals (a function of neural network weights or their updates), not them separately. The users transmit simultaneously in over-the-air FL [150], which exploits the additive nature of the wireless channel. This, alongside the huge number of potential users in envisioned FL applications, makes URA a suitable framework for the physical layer [151, 152]. Challenges of URA in FL: Note that in distributed function computation applications, such as FL, where the receiver is not interested in the transmitted messages themselves but only in a function of them (e.g., their summation in FL), PUPE is not a suitable metric for the URA implementation, as it is defined on the individual messages. So, new metrics (even higher-level ML-based ones such as accuracy) need to be employed to better capture the system performance.
VI-D Integrated Sensing and Communication in URA

### v3

@@ V-F Existing URA Solutions on Alternative Setups
Most URA schemes in the literature consider direct links between all the users and the BS; however, in certain environments, the direct link between some users and the BS may be blocked or significantly attenuated. In these cases, one solution is to employ RIS, a promising technology that can furnish the URA system with high spectral efficiency and energy savings. Specifically, it has been shown that a passive RIS equipped with many low-cost passive elements, which can intelligently tune the phase-shift of the incident electromagnetic waves, and reflect them in a desired direction without any amplification, can improve the efficiency of a URA system by enabling line-of-sight paths between the transmitters and the receiver in problematic environments with many blocking obstacles [134, 135]. More specifically, the RIS-based schemes in [136, 137] operate in two phases. In the first phase, known as the RIS configuration phase, the BS detects the active pilots and estimates their respective channel coefficients, using which the BS suitably selects the phase shifts of the RIS elements. In the second phase, known as the data phase, the transmitted messages of active users are decoded using SIC-based decoders. The RIS-based URA solution in [136] outperforms the one in [137], with a power savings of up to 10 dB. This improvement is due to the employment of a slotted transmission structure, which reduces the interference, as well as a more suitable metric in the RIS design in [136], where the overall phase-shift matrix is obtained by minimizing the decoding error probability, in contrast to maximizing the minimum channel gain among active users for the RIS design in [137]. The authors enhance the RIS design metrics presented in [136] and further improve upon them in [138], obtaining bounds on the system performance.
@@
Conventionally, in MIMO URA systems, ideal hardware units are assumed at both the BS and the UE sides. However, due to the massive number of users and antenna elements, the hardware needs to be inexpensive, and these elements are prone to hardware impairments (HWI) such as I/Q imbalance, phase noise, and quantization errors. The problem of URA with HWI is considered in [139], where the solution developed in [14] is adapted to the presence of HWI.
@@
While most URA works consider transmission over i.i.d. Rayleigh block-fading channels, some investigate other channel models. A channel model with one single dominant path between a user and the BS is considered in [140], and an approach similar to the one in [124] is proposed, exploiting the user location geometry by linking the transmission features such as the employed preamble or interleaver to the angle of arrival of the user. The dynamics of the user states and channels are taken into account in a slotted structure in [141], namely, the users can start or end their transmission in any slot, and they can be active in a subset of the slots, and the channel coefficients change from one slot to another. The authors design a two-step decoding algorithm with an inner AMP and an outer variational message passing. Type-based MA, where each device observes the state of a process and transmits it to a common receiver, is investigated in an unsourced setup in [142], where the receiver’s goal is to estimate the states and their multiplicity. An achievability bound quantifying the performance of type-based MA is derived in [143]. Moreover, the authors in [144] extend the framework in [142] to the case of fading channels, where a cell-free massive MIMO setup is considered, and a solution exploiting location-based codebook partitioning is proposed. The problem of user identification and authentication is applied to URA in [145] with the aid of a message authentication code. Moreover, the authors in [146] propose a URA scheme that combines physical and medium access layers where the medium access layer slotting carries information. The results show that carrying information by the medium access control layer improves the system performance. A Rician fading channel is considered in [147] through an application of coded compressed sensing. A correlated fading scenario is tackled in [148]. Frequency-selective fading is examined in the context of MIMO URA in [149] and [150]. The authors in [149] propose a tensor-based modulation scheme, while a slotted scheme employing OFDM is studied in [150].
@@
The asynchronous setup is also receiving increasing attention due to its easier implementation in hardware. See [151, 153, 152] for proposed solutions in the MIMO URA setting. A slotted concatenated coding scheme is proposed in [151] based on SPARCs and tree codes in an OFDM system, where a variation of the ML technique is employed for decoding of the SPARC. The authors in [152] tackle both timing and frequency offsets and propose an asynchronous URA scheme that employs SPARCs, LDPC codes, and multistage orthogonal pilots for transmission. This scheme also incorporates a graph-based channel reconstruction and collision resolution (GB-CR) algorithm for channel estimation and a modified message passing algorithm for decoding. Symbol asynchronicity in URA is investigated in [153].
@@
Furthermore, some works have investigated URA in other communication setups such as cell-free MIMO [154, 155, 156, 157, 158, 159] and erasure [160, 161] channels. For instance, a cell-free setup is considered in [154] where the coverage area is divided into multiple regions, each served by a receiver unit. Users in each region employ a common codebook, distinct from those in other regions, to transmit their messages. The signals picked up at the receiver units are then jointly processed in a central unit. Joint processing is also considered in [156] and [157], where a combination of channel coding and fixed sparse coding is employed in the former, while an uncoupled CCS scheme utilizing coarse location information for trellis stitching is proposed in the latter. On the other hand, Level 2 cooperation, where only the symbol estimates are passed from the access points to the central processing unit, is considered in [158] and [159]. Both schemes are pilot-based, and [158] employs the random spreading approach in the data part, while ODMA is utilized in [159].
@@
The common assumption in the URA literature is that all users have the same number of bits to transmit or require the same quality of service. However, this may be impractical in situations where the devices have different payload requirements or power budgets, especially considering their massive number. This problem, known as multi-class URA, is addressed in [162, 163] within the coded compressed sensing framework. More specifically, a scenario is considered in which devices are clustered into two classes with different SNR levels and payload requirements. In the cluster with higher power, devices transmit using a two-layer superposition modulation, while those in the other cluster transmit with the same base constellation as in the first cluster. To recover packets at the receiver, a CCS-based approach is employed, where an outer code is used to stitch message fragments together across time and layers.
We provide a summary of different URA setups and the corresponding solution approaches in Fig. 20.
Figure 20: A summary of different solutions in the URA literature.
VI Conclusions
In this survey, we provided a comprehensive review of the URA literature, with an emphasis on the state-of-the-art. We grouped different URA solutions according to their assumptions on the underlying wireless channels, and explained the representative works in each group in more detail. Moreover, we compared the works in the literature both in terms of setup and performance to gain a clearer understanding of their pros and cons. Across the three major channel models, our survey demonstrates a consistent evolution of URA research: from foundational achievability bounds on the GMAC, to practical yet computationally constrained designs in single-antenna fading channels, and finally to scalable, high-performing solutions enabled by MIMO. This progression reflects the natural trajectory of URA from theory toward real-world adoption in 6G and IoT systems. The comparative analyses highlight that future URA solutions must jointly optimize energy efficiency, interference management, and computational complexity, while also harnessing spatial diversity and structural coding. Taken together, these insights position URA not only as a theoretical construct but as a practical enabler of massive connectivity, capable of meeting the stringent requirements of next-generation networks.
VI-A Lessons Learned
The key lessons and insights learned from existing URA solutions can be summarized as follows.
  • ODMA is a promising URA solution. State-of-the-art URA solutions provide a balance between energy efficiency, scalability, and low-complexity decoding. Among these solutions, ODMA-based schemes excel by offering an effective trade-off between performance, sparsity, and computational complexity. The time-domain sparsity of ODMA-based solutions significantly reduces user interference, making multiuser detection and decoding more manageable even under high user loads, resulting in outstanding performance in both AWGN and fading scenarios, particularly when combined with power diversity. The structured sparsity introduced by ODMA is especially advantageous in massive MIMO systems, where spatial diversity can further enhance separation and decoding accuracy. Additionally, ODMA-based solutions are promising for asynchronous URA, as the ODMA transmission patterns can be utilized instead of commonly employed concatenated preambles to estimate the start times of packets necessary for asynchronous detection, thereby increasing system efficiency. Overall, ODMA-based schemes are promising candidates for scalable, practical, and high-performance solutions for anticipated URA applications, such as massive IoT.
  • Slotted transmission is a simple yet effective interference management technique in URA. Interference management is one of the major challenges in URA. Among various interference management approaches in the URA literature, ODMA and slotted transmission are promising yet straightforward techniques that introduce structured sparsity by limiting the transmission of user packets to a small portion of the frame, thereby reducing interference. More specifically, by dividing the transmission frame into multiple slots and allowing each user to randomly select one or more slots for transmission, slotting-based solutions decrease the number of simultaneously active users per slot. This localized activity reduces multiuser interference, thereby enabling more accurate signal recovery and effective use of decoding techniques such as SIC. Furthermore, slotted transmission simplifies the decoding process by confining computations to smaller subsets of users, thus decreasing the overall computational complexity compared to full-frame schemes. As a result, slotting serves as a key design element for practical and low-complexity URA implementations.
  • Pilot design and proper channel estimation are critical in URA scenarios over fading channels. In fading environments, reliable channel estimation is essential for accurate user separation and decoding. Since assigning dedicated pilots to each user is infeasible in URA due to scalability constraints, pilot-based solutions often require users to draw pilot sequences from a common, non-orthogonal codebook. This introduces the risk of pilot collisions, which can lead to severe interference and decoding failures. Therefore, minimizing the probability of pilot collisions and optimizing pilot properties, such as length, power, and structure, are crucial for effective activity detection, channel estimation, and interference management. Note that existing pilot-free URA schemes eliminate pilot overhead, but often result in higher complexity and less reliable performance compared to their pilot-based counterparts.
  • Massive MIMO is a game changer in the realization of URA schemes. Provided with a large number of antennas, the URA receiver can better manage interference and enhance channel estimation accuracy, even under heavy user loads. In particular, massive MIMO enables advanced decoding strategies, such as spatial separation of users and improved compressed sensing techniques, that would be infeasible for a receiver with a limited number of antennas. These capabilities position massive MIMO as a foundational enabler for highly efficient, scalable communication in a URA setup.
  • URA schemes are sensitive to various system parameters. Certain system parameters can significantly affect the performance, computational complexity, and scalability in URA solutions. Therefore, key parameters, such as the number of active users, frame length, number of slots, pilot sequence length, and power allocation, must be carefully optimized to balance trade-offs between interference mitigation, decoding accuracy, and computational efficiency. Improper tuning of these parameters can lead to severe performance degradation; for instance, too many pilot sequences increase the computational complexity of activity detection, while insufficient pilot diversity raises the risk of collisions. The interdependence of these factors makes robust URA design a delicate task, requiring adaptive parameter configuration based on traffic patterns and deployment scenarios to ensure reliable communication.
VI-B Future Research Directions
@@
As evident from this survey, numerous works on URA have been conducted in recent years. However, there are still issues, some of which are highly critical, that require further investigation. We will next outline some of these issues and potential methods to address them, and provide several promising future research directions.
VI-B1 Asynchronous URA
@@
In URA, the common assumption is that the users transmit simultaneously, or more precisely, their signals are received at the same time at the receiver (e.g., symbol synchronicity). However, due to the nature of users and devices in envisioned URA applications, such as massive sensor networks, this transmission coordination is very challenging. This makes asynchronous URA an important and practically relevant research direction to investigate. Note that even if the transmission synchronization is realized, the synchronous reception cannot be guaranteed. Due to the massive number of potential devices with diverse distances to the receiver, the difference in reception times between many users can be larger than a time slot, which makes the setup asynchronous. Interestingly, it has been recently shown that if asynchronicity is dealt with properly, it can even lead to performance improvements compared to the synchronous case [164].
@@
There have been some recent attempts in the literature to address asynchronous URA [100, 99, 103, 151, 152, 51]. Assuming a bounded time horizon enables these works to convert the problem to one similar to the synchronous scenario, e.g., by employing a limited decoding time frame. For instance, in [100], the authors employ a frequency-domain solution to convert the asynchronous problem to a phase-shift estimation one. Then, conventional synchronous strategies such as TIN-SIC are utilized to decode the active users. On the other hand, [81] and [87] consider a fully asynchronous (i.e., with an unbounded time horizon) setup for URA over GMAC, and a window-based decoding approach.
@@
The primary URA-specific challenge in a fully asynchronous setup for URA over fading channels is the increased computational complexity, which is significantly higher than that of a conventional (non-URA) setup due to the higher number of required preambles. More specifically, in an asynchronous GMAC setup, the preambles only serve to identify the start of the packets at the receiver, so even a very small preamble pool would suffice; while in an asynchronous fading setup, they are also needed to estimate the channel coefficients, which are then used for user separation. In a conventional fading setup, the receiver needs to search only over the assigned preambles, whose number equals the number of active users. However, in a URA setup, due to the lack of coordination between the users and the receiver, the number of available preambles must be significantly higher than in the conventional setup; hence, a much higher computational complexity is required for user activity detection and channel estimation. This computational complexity substantially increases in the asynchronous setup with a sliding window receiver, as the receiver shifts and correlates all possible preambles through the window to find the start of the packets as well as the user channel estimates, repeating this process for each shift of the sliding window. This significantly increases the overall computational complexity, rendering the solution practically infeasible. One way to reduce the complexity is to increase the moving steps of the sliding window, at the cost of an overall performance loss. More sophisticated schemes are needed to alleviate this issue.
VI-B2 Robust URA
@@
Like any other communication system, the performance of a URA system is affected by the presence of imperfections in the assumed models or employed communication blocks. For instance, the presence of hardware impairments, e.g., due to non-linearities of the power amplifiers or phase errors in phase-shifters, is highly probable in many practical URA implementations due to cheap and low-precision devices. So far, this challenge is addressed only in [139], where a robust solution is presented for the case of residual hardware impairments in which the remaining hardware impairments after compensation are statistically modeled as additive and multiplicative noises. The analysis of different hardware impairments on the performance of existing URA solutions and providing robust solutions to mitigate them is an interesting research direction.
@@
Examining the URA literature, we can observe that URA schemes are more sensitive to the choice of design parameters than conventional ones. That is, for the URA schemes to achieve acceptable performance, their design parameters must be optimized whenever a system hyperparameter, such as the active user load or the target PUPE, changes. This makes their standardization and practical employment challenging. Therefore, modifications are needed to make the URA solutions robust against changes in system hyperparameters. One approach is to employ stochastic optimization techniques, such as optimizing the expected value of the cost metric (e.g., PUPE) over the statistics of the intended system hyperparameters, to obtain a set of design parameters that perform well across a wide range of system hyperparameters.
VI-B3 Machine Learning (ML) for and with URA
@@
Current solutions look at the URA problem in a single shot. However, the transmission is recurrent and spans several frames in many practical applications. In such cases, ML techniques, such as reinforcement learning, can be employed to adapt to the environment and learn suitable transmission strategies. ML techniques can also be used to simplify certain blocks in the decoding process, such as user activity detection (similarly to the deep learning-based solutions in [165]) and channel estimation. For instance, in [166], the authors demonstrate that employing a dictionary learning-based solution eliminates the need for pilots for channel estimation in URA scenarios over fading channels while maintaining an acceptable performance.
Apart from general challenges in implementing ML solutions in communication systems, such as energy and computational constraints and compatibility with legacy systems, there are some that are specific or more pronounced in a URA setup, which is mostly due to the massive number of potential users (up to the order of millions) and their sporadic activity. Training a neural network with an input of very large size (due to the massive number of potential users) with non-stationary characteristics (due to the sporadic activity) is a very challenging task.
URA can also be employed to enable distributed learning among wireless agents. In federated learning (FL), the goal of the parameter server, which acts as the orchestrator, is to aggregate local model updates from many clients at each iteration, and update the global model by averaging these models [167]. When clients are co-located, and the parameter server is a wireless access point, model updates have to be communicated over a shared wireless channel. In conventional methods, this requires sharing the available channel resources among the active clients; however, in FL, the parameter server is not interested in individual updates. Therefore, in [168], over-the-air computing (OAC) was proposed to exploit the signal superposition property of wireless signals. This would allow the parameter server to directly receive the summation of the model updates.
However, such an analog over-the-air computing scheme would require a certain level of coordination and synchronization among the clients. An alternative digital OAC scheme is proposed in [169] using URA. In OAC with URA, each client quantizes its model update, and sends the corresponding codeword. Both the quantization and channel codebooks are shared among all the clients. Differently from conventional URA, which relies on the fact that active users are likely to transmit independent codewords, and the receiver tries to identify only the transmitted codewords by the active clients, here the users are more likely to transmit the same codeword as their updates tend to converge as the training process progresses. Moreover, the receiver wants to identify not only the active codewords, but also how many clients transmit each codeword. This approach can be extended to other functions, and the identification of the optimal quantization and communication codebooks is an open research direction.
VI-B4 Semantic and Task-Oriented URA
Conventional URA frameworks are designed under the assumption that all transmitted bits are equally important, focusing primarily on reliable message recovery. However, in many IoT and machine-type communication scenarios, the goal is not to recover raw data streams but rather to support specific tasks such as anomaly detection, control signaling, or status monitoring. In such cases, transmitting semantic or task-oriented information, i.e., only the most relevant features of the data, can drastically reduce the required payload and improve spectral efficiency. Future research may investigate how semantic encoding techniques can be combined with URA to enable resource-efficient access, for example, by allowing multiple devices reporting correlated measurements to share compressed semantic representations. Such integration would align URA more closely with emerging 6G paradigms, where communication systems are expected to deliver knowledge and actionable insights rather than just data.
VI-B5 ISAC in URA